# Data-free Asymptotics-Informed Operator Networks for Singularly Perturbed PDEs

### v1

@@ 2.3 Convergence of eFEONet
In this section, we discuss the convergence result for eFEONet, providing a theoretical foundation for the proposed approach. Although the convergence analysis of eFEONet does not differ substantially from that of the original FEONet lee2024finite, we include a concise exposition in this section in order to introduce a coherent framework for the theoretical analysis of methods of this type. We let an external forcing term $f$ as the input of neural networks, that is parametrized by $\boldsymbol{\omega}$ in the probability space $(\Omega,\mathcal{F},\mathbb{P})$. We shall interpret $f(\mathbf{x};\boldsymbol{\omega})$ as a bivariate function defined on $D\times\Omega$. Moreover, we will assume that
@@
Next, we define the class of feed-forward neural networks as $\mathcal{N}_{n}$, where the subscript $n$ denotes the network architecture. We assume that $\mathcal{N}_{n_{2}}$ is more expressive than $\mathcal{N}_{n_{1}}$ when $n_{1}\leq n_{2}$. For instance, $n$ could represent the number of layers with bounded width, or the number of neurons when the number of layers is fixed. Neural networks are known to be an appropriate choice for nonlinear approximation, supported by the universal approximation theorem (see, for example, (UA_1; UA_2; UA_3; terry_lion)). Now for a neural-network approximation of $\alpha^{*}$, we mean that $\widehat{\alpha}(n):\Omega\rightarrow\mathbb{R}^{N(h)+1}$, which solves the following minimization problem
@@
Here $\mathcal{L}^{M}$ is the empirical risk, which is the Monte–Carlo integration of the population risk $\mathcal{L}(\alpha)$:
@@
The first error arises from the finite element approximation, which we assume to be negligible when $h>0$ is sufficiently small. The error analysis for the first term (I) is well investigated in the previous literature on singular perturbation analysis. For example, in (BLA_1), the following error estimate was derived for the enriched FEM (2.2):
@@
where $C>0$ is a constant independent of $h$ and $\varepsilon$. This result is especially highlighted as it provides a satisfactory convergence result even in the under-resolved case for $h>\varepsilon$. A mathematical analysis of this problem constitutes an independent topic traditionally addressed within classical numerical analysis; accordingly, we shall not examine it in detail in the present work. More general results can be found in various papers, e.g., from (BLA_2; BLA_3). For the analysis of eFEONet, based on the estimate (19), we can reduce this error to any desired level by selecting a suitable $h>0$. Therefore, we assume that $h$ has been chosen so that the finite element approximation error is small enough. The second error, known as the approximation error, occurs when we use a class of neural networks to approximate the target (finite element) coefficients. The third error, often referred to as the generalization error, measures how well our approximation performs on unseen data. Our focus will be on proving that, with fixed $h>0$ and $\varepsilon>0$, as the index $n\in\mathbb{N}$ for neural network architectures becomes larger and the number of input samples $M\in\mathbb{N}$ increases, our approximate solution $\widehat{u}^{\rm en}_{\varepsilon,h,n,M}$ converges to the finite element solution $u^{\rm en}_{\varepsilon,h}$ which is assumed to be the true solution here.
@@
As mentioned earlier, this theorem can be shown directly from the proof presented in the original paper concerning FEONet lee2024finite. The only difference lies in the singular perturbation analysis, which governs (19), while the approximation and generalization errors retain a similar structure for a fixed $\varepsilon>0$. Unlike original FEONet, however, the associated constants in our setting depend implicitly on the perturbation parameter $\varepsilon$. It is noteworthy that the convergence in Theorem 2.3 is not uniform with respect to $h\rightarrow 0$. Indeed, this issue aligns precisely with the main theme of the reference (feonet_anal), where the authors rigorously demonstrated that both the approximation error and the generalization error depend on the condition number $\kappa(A)$ of the finite element matrix $A$, which can typically be estimated as a scale $\kappa(A)\sim h^{-2}$. This means that as $h$ becomes smaller, both the approximation and generalization errors may increase due to this adverse dependence. While this analysis was originally developed in the context of FEONet, it applies directly to eFEONet as well, since in eFEONet we solve equations with a fixed small $\varepsilon$. More precisely, the only part of the analysis in (feonet_anal) where $\varepsilon$ could potentially affect the results is in the condition number estimates. If we explicitly characterize the dependency on $\varepsilon$ in these equations, then we can likewise make the $\varepsilon$ -dependence explicit in the final error estimate. In doing so, we can obtain a complete error analysis for eFEONet that incorporates both singular perturbation asymptotic analysis and the general framework from (feonet_anal), which will be addressed in the forthcoming paper.
@@ 3 Numerical Experiments
In this section, we evaluate the performance of eFEONet on three distinct types of singularly perturbed differential equations, including both ordinary and partial differential equations. For ordinary differential equations (ODEs), we examine scenarios with and without turning points, highlighting eFEONet’s adaptability to varying problems. In case of PDEs, we assess the performance and robustness of eFEONet on two-dimensional problems defined on square domain. Furthermore, we conduct a comparison of the experimental results with those obtained using FNO (kovachki2021neural) and ComFNO (SP03), a neural operator model specifically designed to address the challenges of singularly perturbed differential equations.
The high-precision numerical solutions are denoted as $u_{\varepsilon}$, while the predictions are represented as $\widehat{u}_{\varepsilon}$. The training dataset consists of 900 load vectors generated from independently sampled functions $f$, with inputs discretized at a resolution of 201 for both 1D and 2D cases.
In order to train neural networks, we need to generate random external forcing functions. Inspired by (bar2019learning), we created a random signal $f(\mathbf{x};\boldsymbol{\omega})$ as a linear combination of sine functions and cosine functions. More precisely, we set
$f(x)=m_{0}\sin(n_{0}x)+m_{1}\cos(n_{1}x)$ (21)
for 1D cases and
$f(x,y)=m_{0}\sin(n_{0}x+n_{1}y)+m_{1}\cos(n_{2}x+n_{3}y)$ (22)
for 2D cases where $m_{i}$ for $i=1,2$ and $n_{j}$ for $j=0,1,2,3$ are drawn independently from the uniform distributions. It is worth noting that even when considering different random input functions, such as those generated by Gaussian random fields, we consistently observe similar results. This robustness indicates the reliability and stability of the eFEONet approach across various input scenarios.
High-precision numerical solutions on the Shishkin mesh (see, e.g., (SP03)) are used to compute the corresponding outputs $u_{\varepsilon}$, which serve as the ground truth during training. Additionally, for all ODE experiments, the input-output resolution is set to 201, ensuring consistency across the comparative evaluations of FNO, ComFNO, and our method. In 2D PDE experiments, the resolution is fixed at 51 for $\varepsilon=10^{-3}$ and for $\varepsilon=10^{-4}$ in the rectangular domain.
The overall experimental results are summarized in Table 3. We compare our method with several other operator-learning approaches, and we also vary the number of training samples to assess the dependence on training data. As shown in the table, our method substantially outperforms competing methods that require training data, even when our approach is trained without using any training data.
@@ 3.1 Experimental Setup and Implementation Details
In this section, we describe the experimental setup used to ensure fair and reproducible comparisons across all methods. We employ a neural network architecture consisting of six convolutional layers with Swish activation functions, followed by a fully connected layer. For 1D problems, the convolutional layers are implemented with standard one-dimensional convolutions (Conv1D), whereas for 2D problems they are implemented with two-dimensional convolutions (Conv2D). The eFEONet was trained using the L-BFGS optimizer with the following hyperparameters:
@@
All experiments were conducted on an Intel Xeon Gold 6226R CPU and an NVIDIA RTX A6000 GPU (48GB). For the 1D problems, the training dataset for FNO and ComFNO includes $900\times 201$ tuples $(f,u)$, while the 2D scenarios encompass $900\times 51\times 51$ tuples $(f,u)$ as described in the paper (SP03). In all experiments, we employed the mean squared error (MSE) loss. For FNO and ComFNO, we used the Adam optimizer for all minimization problems, accompanied by the consistent utilization of the GELU activation function. Further details concerning the remaining parameters for our result can be found in Table 1 and Table 2.
3.2 Ordinary Differential Equations with Boundary Layer
@@
$-\varepsilon u^{\prime\prime}_{\varepsilon}+(x+1)u^{\prime}_{\varepsilon}$ $=f(x),x\in(0,1),$ (23)
@@
The performance of our method for $\varepsilon=10^{-5}$ with 100 test function $f$ samples is shown in Figure 4. The figure on the left shows the input function $f$, the middle figure shows the ground truth corresponding to the 100 test $f$ samples, and the figure on the right shows the residuals produced by our method for these 100 $f$ samples.
Figure 4: Visualization of 100 input functions $f$ (left), corresponding reference solutions (middle), and error plots (right) for boundary layer problem (23) with $\varepsilon=10^{-5},\text{input-output resolution}=51$.
@@ 3.3 Ordinary Differential Equations with Interior Layer
with the corrector function $\phi_{0}(x)={\rm erf}(\sqrt{1/(2\varepsilon)}x)$. As shown in the third column of Table 3, eFEONet achieves better accuracy than both FNO and ComFNO, with a larger performance gap emerging as the number of training samples decreases. This highlights the robustness of eFEONet in data-scarce scenarios. Table 5 shows the relative $L^{2}$ test errors for FNO, ComFNO, and eFEONet across different values of $\varepsilon$ for ODEs with interior layers. The results demonstrate that eFEONet consistently achieves superior accuracy compared to FNO and ComFNO, even in the absence of training data. Notably, as $\varepsilon$ decreases, the performance gap between eFEONet and the benchmark models significantly widens, indicating eFEONet’s ability to accurately capture sharp interior layers.
Figure 6 compares the predicted solutions $\widehat{u}_{\varepsilon}$ for two test samples using FNO, ComFNO, and eFEONet with $\varepsilon=10^{-8}$. Notably, eFEONet demonstrates superior accuracy, particularly around the singular region near the turning point at $x=0$. This result underscores the capability of eFEONet to effectively handle the challenges posed by singularities and turning points in differential equations, delivering reliable predictions even in complex scenarios. Figure 7 shows the input function $f$ (left), the ground truth corresponding to the 100 test $f$ samples (middle), and the residuals produced by our method for these 100 $f$ samples (right).
@@
Figure 7: Visualization of 100 input functions $f$ (left), corresponding reference solutions (middle), and error plots (right) for interior layer problem (24) with $\varepsilon=10^{-5},\text{input-output resolution}=51$.
@@ 3.4 Partial Differential Equations on Square
For a boundary-value problem of an elliptic PDE in the spatial domain $D=[0,1]^{2}$, we consider
@@
$u_{\varepsilon}(x,y)$ $=0\,\,\text{on}~\partial D,$
where the solution exhibits a boundary layer along the edge at $x=0$ and $y=0$, as illustrated in Figure 8. For this PDE problem, the asymptotic expansion of $u(x,y)$ is formulated as:
@@
As shown in the fourth column of Table 3, the accuracy gap between eFEONet and benchmark models becomes even more pronounced for this problem. This highlights the capability of eFEONet to effectively resolve boundary layers in complex spatial domains. Furthermore, as seen in Figure 9, ComFNO shows large errors, whereas eFEONet achieves consistently low errors across the entire domain, demonstrating its robustness and superior accuracy in handling such challenging scenarios.
Figure 9: Comparison of the reference solution $u_{\varepsilon}(x,y)$ (left) and the predicted solutions $\widehat{u}_{\varepsilon}$ along the diagonal $y=x$ for ComFNO and eFEONet (middle and right) with $\varepsilon=10^{-4}$. The results highlight the superior accuracy of eFEONet in capturing sharp boundary layers along $x=0$, whereas ComFNO exhibits noticeable errors near the boundary regions.
The solution exhibits a boundary layer along $x=0$ and $y=0$, with a corner layer forming at $(0,0)$, as illustrated in Figure 8. To solve the equation (25), we employ both FNO, ComFNO, and eFEONet. Residuals for 100 randomly chosen $f$ sample with $\varepsilon=10^{-4}$ are presented in Figure 10.
Figure 10: Visualization of error plots from FNO for 100 test function(left), ComFNO(middle) and eFEONet(right) for Section 4.3 with $\varepsilon=10^{-4},\text{input-output resolution}=51$.

### v2

@@ 2.3 Convergence of eFEONet
In this section, we discuss the convergence result for eFEONet, providing a theoretical foundation for the proposed approach. Although the convergence analysis of eFEONet does not differ substantially from that of the original FEONet Lee et al. (2025), we include a concise exposition in this section in order to introduce a coherent framework for the theoretical analysis of methods of this type. We let an external forcing term $f$ as the input of neural networks, that is parametrized by $\boldsymbol{\omega}$ in the probability space $(\Omega,\mathcal{F},\mathbb{P})$. We shall interpret $f(\mathbf{x};\boldsymbol{\omega})$ as a bivariate function defined on $D\times\Omega$. Moreover, we will assume that
@@
Next, we define the class of feed-forward neural networks as $\mathcal{N}_{n}$, where the subscript $n$ denotes the network architecture. We assume that $\mathcal{N}_{n_{2}}$ is more expressive than $\mathcal{N}_{n_{1}}$ when $n_{1}\leq n_{2}$. For instance, $n$ could represent the number of layers with bounded width, or the number of neurons when the number of layers is fixed. Neural networks are known to be an appropriate choice for nonlinear approximation, supported by the universal approximation theorem (see, for example, (Cybenko, 1989; Hornik, 1991; Pinkus, 1999; Kidger and Lyons, 2020)). Now for a neural-network approximation of $\alpha^{*}$, we mean that $\widehat{\alpha}(n):\Omega\rightarrow\mathbb{R}^{N(h)+1}$, which solves the following minimization problem
@@
Here $\mathcal{L}^{M}$ is the empirical risk, which is the Monte Carlo integration of the population risk $\mathcal{L}(\alpha)$:
@@
The first error arises from the finite element approximation, which we assume to be negligible when $h>0$ is sufficiently small. The error analysis for the first term is well investigated in the previous literature on singular perturbation analysis. For example, in (Cheng and Temam, 2002), the following error estimate was derived for the enriched FEM (2.2):
@@
where $C>0$ is a constant independent of $h$ and $\varepsilon$. This result is especially highlighted as it provides a satisfactory convergence result even in the under-resolved case for $h>\varepsilon$. A mathematical analysis of this problem constitutes an independent topic traditionally addressed within classical numerical analysis; accordingly, we shall not examine it in detail in the present work. More general results can be found in various papers, e.g., from (Jung, 2005; Gie et al., 2018). For the analysis of eFEONet, based on the estimate (19), we can reduce this error to any desired level by selecting a suitable $h>0$. Therefore, we assume that $h$ has been chosen so that the finite element approximation error is small enough. The second error, known as the approximation error, occurs when we use a class of neural networks to approximate the target (finite element) coefficients. The third error, often referred to as the generalization error, measures how well our approximation performs on unseen data. Our focus will be on proving that, with fixed $h>0$ and $\varepsilon>0$, as the index $n\in\mathbb{N}$ for neural network architectures becomes larger and the number of input samples $M\in\mathbb{N}$ increases, our approximate solution $\widehat{u}^{\rm en}_{\varepsilon,h,n,M}$ converges to the finite element solution $u^{\rm en}_{\varepsilon,h}$ which is assumed to be the true solution here.
@@
As mentioned earlier, this theorem can be shown directly from the proof presented in the original paper concerning FEONet Lee et al. (2025). The only difference lies in the singular perturbation analysis, which governs (19), while the approximation and generalization errors retain a similar structure for a fixed $\varepsilon>0$. Unlike the original FEONet, however, the associated constants in our setting depend implicitly on the perturbation parameter $\varepsilon$. It is noteworthy that the convergence in Theorem 2.3 is not uniform with respect to $h\rightarrow 0$. Indeed, this issue aligns precisely with the main theme of the reference (Hong et al., 2025), where the authors rigorously demonstrated that both the approximation error and the generalization error depend on the condition number $\kappa(A)$ of the finite element matrix $A$, which can typically be estimated as a scale $\kappa(A)\sim h^{-2}$. This means that as $h$ becomes smaller, both the approximation and generalization errors may increase due to this adverse dependence. While this analysis was originally developed in the context of FEONet, it applies directly to eFEONet as well, since in eFEONet we solve equations with a fixed small $\varepsilon$. More precisely, the only part of the analysis in (Hong et al., 2025) where $\varepsilon$ could potentially affect the results is in the condition number estimates. If we explicitly characterize the dependency on $\varepsilon$ in these equations, then we can likewise make the $\varepsilon$ -dependence explicit in the final error estimate. In doing so, we can obtain a complete error analysis for eFEONet that incorporates both singular perturbation asymptotic analysis and the general framework from (Hong et al., 2025), which will be addressed in the forthcoming paper.
@@ 3 Numerical Experiments
In this section, we evaluate the performance of eFEONet on three distinct types of singularly perturbed differential equations, including both ordinary and partial differential equations. For ordinary differential equations (ODEs), we examine scenarios with and without turning points, highlighting eFEONet’s adaptability to varying problems. In the case of PDEs, we assess the performance and robustness of eFEONet on two-dimensional problems defined on a square domain. In all experiments considered in this section, eFEONet employs corrector bases derived via asymptotic analysis; the derivations are provided in the appendix. Furthermore, we conduct a comparison of the experimental results with those obtained using FNO (Kovachki et al., 2021) and DeepONet (Lu et al., 2021a), the well-known operator-learning methods, and ComFNO (Li et al., 2024), the neural operator model specifically designed to address the challenges of singularly perturbed differential equations. The overall experimental results are summarized in Table 3. We vary the number of training samples to assess the dependence on training data.
@@ 3.1 Experimental Setup and Implementation Details
In this section, we describe the experimental setup used to ensure fair and reproducible comparisons across all methods. In order to train neural networks, we need to generate random external forcing functions. Inspired by (Bar-Sinai et al., 2019), we created a random signal $f(\mathbf{x};\boldsymbol{\omega})$ as a linear combination of sine functions and cosine functions. More precisely, we set
$f(x)=m_{0}\sin(n_{0}x)+m_{1}\cos(n_{1}x),$ (21)
for 1D cases and
$f(x,y)=m_{0}\sin(n_{0}x+n_{1}y)+m_{1}\cos(n_{2}x+n_{3}y),$ (22)
for 2D cases where $m_{i}$ and $n_{j}$ are sampled from the standard normal distribution. It is worth noting that even when considering different random input functions, such as those generated by Gaussian random fields, we consistently observe similar results. This robustness indicates the reliability and stability of the eFEONet approach across various input scenarios.
The high-precision numerical solutions are denoted as $u_{\varepsilon}$, while the predictions are represented as $\widehat{u}_{\varepsilon}$. High-precision numerical solutions on the Shishkin mesh (see, e.g., (Li et al., 2024)) are used to compute the corresponding outputs $u_{\varepsilon}$, which serve as the ground truth during training. The training dataset comprises 900 load vectors derived from functions $f$ sampled independently. These functions are discretized using a resolution of 201 points for 1D cases and a $51\times 51$ grid for 2D cases. Additionally, for all ODE experiments, the input-output resolution is set to 201, ensuring consistency across the comparative evaluations of FNO, ComFNO, DeepONet, and our method. In 2D PDE experiments, the resolution is fixed at 51 for $\varepsilon=10^{-3}$ and for $\varepsilon=10^{-4}$ in the rectangular domain.
We employ a neural network architecture consisting of six convolutional layers with Swish activation functions, followed by a fully connected layer. For 1D problems, the convolutional layers are implemented with standard one-dimensional convolutions (Conv1D), whereas for 2D problems they are implemented with two-dimensional convolutions (Conv2D). The eFEONet was trained using the L-BFGS optimizer with the following hyperparameters:
@@
All experiments were conducted on an Intel Xeon Gold 6226R CPU and an NVIDIA RTX A6000 GPU (48GB). For the 1D problems, the training dataset for FNO, ComFNO, and DeepONet includes $900\times 201$ tuples $(f,u)$, while the 2D scenarios encompass $900\times 51\times 51$ tuples $(f,u)$ as described in the paper (Li et al., 2024). In all experiments, we employed the mean squared error (MSE) loss. For FNO and ComFNO, we used the Adam optimizer for all minimization problems, accompanied by the consistent utilization of the GELU activation function. Further details concerning the remaining parameters for our result can be found in Table 1 and Table 2. For DeepONet, we employed the standard branch-trunk architecture. Both the branch and trunk networks consist of three hidden layers with 100 neurons and tanh activation functions. The model was optimized using full batch and the Adam optimizer with a learning rate of $10^{-4}$ for $100,000$ training epochs.
@@
3.2 Ordinary Differential Equations with Boundary Layer
@@
$-\varepsilon u^{\prime\prime}_{\varepsilon}+(x+1)u^{\prime}_{\varepsilon}$ $=f(x),\quad x\in(0,1),$ (23)
@@
Figure 5 illustrates the performance of eFEONet for $\varepsilon=10^{-5}$ using 100 test samples of the forcing function $f$ at a resolution of 51. The figure on the left shows the input function $f$, the middle figure shows the ground truth corresponding to the 100 test $f$ samples, and the figure on the right shows the residuals produced by our method for these 100 $f$ samples.
@@
Figure 5: Visualization of 100 input functions $f$ (left), corresponding reference solutions (middle), and error plots (right) for boundary layer problem (23) with $\varepsilon=10^{-5}$.
@@ 3.3 Ordinary Differential Equations with Interior Layer
with the corrector function $\phi_{0}(x)={\rm erf}(\sqrt{1/(2\varepsilon)}x)$ where ${\rm erf}(\cdot)$ is the Gauss error function. As shown in the third column of Table 3, eFEONet achieves better accuracy than both FNO and ComFNO, with a larger performance gap emerging as the number of training samples decreases. This highlights the robustness of eFEONet in data-scarce scenarios. Table 5 shows the relative $L^{2}$ test errors for FNO, ComFNO, and eFEONet across different values of $\varepsilon$ for ODEs with interior layers. The results demonstrate that eFEONet consistently achieves superior accuracy compared to FNO and ComFNO, even in the absence of training data. Notably, as $\varepsilon$ decreases, the performance gap between eFEONet and the benchmark models significantly widens, indicating eFEONet’s ability to accurately capture sharp interior layers.
Figure 6 compares the predicted solutions $\widehat{u}_{\varepsilon}$ for two test samples using FNO, ComFNO, and eFEONet with $\varepsilon=10^{-8}$. Notably, eFEONet demonstrates superior accuracy, particularly around the singular region near the turning point at $x=0$. This result underscores the capability of eFEONet to effectively handle the challenges posed by singularities and turning points in differential equations, delivering reliable predictions even in complex scenarios. Figure 7 shows the input function $f$ (left), the ground truth corresponding to the 100 test $f$ samples (middle), and the residuals produced by our method for these 100 $f$ samples (right) at a resolution of 51.
@@
Figure 7: Visualization of 100 input functions $f$ (left), corresponding reference solutions (middle), and error plots (right) for interior layer problem (24) with $\varepsilon=10^{-5}$.
@@ 3.4 Partial Differential Equations on Square
For a boundary-value problem of an elliptic PDE on the spatial domain $D=[0,1]^{2}$, we consider
@@
$u_{\varepsilon}(x,y)$ $=0\quad\,\qquad\text{on}~\partial D,$
where the solution exhibits a boundary layers along the edges at $x=0$ and $y=0$, as illustrated in Figure 8. For this PDE problem, the asymptotic expansion of $u(x,y)$ can be expressed as:
@@
As shown in the fourth column of Table 3, the accuracy gap between eFEONet and benchmark models becomes even more pronounced for this problem. This highlights the capability of eFEONet to effectively resolve boundary layers in complex spatial domains. Furthermore, Figure 9 illustrates that ComFNO exhibits relatively large errors near the boundary layers, whereas eFEONet maintains uniformly low error levels across the domain. This highlights the robustness of eFEONet in accurately resolving sharp boundary-layer structures. We overlaid the error plots for 100 random inputs of $f$ in Figure 10, calculated with $\varepsilon=10^{-4}$.
Figure 9: Comparison of the reference solution $u_{\varepsilon}(x,y)$ (left) and the predicted solutions $\widehat{u}_{\varepsilon}$ along the diagonal $y=x$ plane for ComFNO and eFEONet (middle and right) with $\varepsilon=10^{-4}$. The results highlight the superior accuracy of eFEONet in capturing sharp boundary layers along $x=0$, whereas ComFNO exhibits noticeable errors near the boundary regions.
Figure 10: Error visualizations for 100 test forcing functions using FNO (left), ComFNO (middle), and eFEONet (right) for the PDE (25) on square with $\varepsilon=10^{-4}$ and input–output resolution $51$.